Question 7 Three particles are fixed on an x axis. Particle 1 of charge q1 is at x =-a and particle 2 of charge q2 is at x = +a. If their net electrostatic force on particle 3 of charge Q is to be zero, what must be the ratio qi/q when particle 3 is at (a) x+0.354a and (b)x+2.08a? (a) Number (b) Number Click if you would like to Show Work for this question: Open Show Work Units Units
Answer #1

(a) q1/(a + 0.354a)^2 = q2/(-a + 0.354a)^2

q1/q2 = (1 + 0.354)^2 /(1 - 0.354)^2

= 4.4

(b) q1/(a + 2.08a) = q2/(-a + 2.08a)^2

q1/q2 = (1 + 2.08)^2 /(-1 + 2.08)^2

= 8.13